9x+6y+81
4x+3y=36
my answer is (9.0)

Answer by tommyt3rd(5050)

9x+6y+81
m2: 4x+3y=36

9x+6y=81
8x+6y=72

subtracting yields
x=9
and so y=0

you are correct :)